# Algebraic Bethe ansatz approach to the correlation functions of the one-dimensional bosons with attraction

### v1

@@ 1 Introduction
In this paper, we apply the ABA to the study of correlation functions of a one-dimensional Bose gas with attractive interaction. The Hamiltonian of this model is given by
@@
Parameter $c$ in (1.1) is a coupling constant. The case of attraction corresponds to $c<0$. Therefore, in what follows we set $c=-\varkappa$ and consider $\varkappa>0$.
The spectral problem for this model was solved in [37, 38]. The ground state energy and the wave function in the attractive case were studied in [39, 40]. A formulation of the model on the ABA language was given in [41, 42]. A series of multiple integrals for correlation functions of this model was derived in [43, 44]. In [29], a representation for the correlation function was obtained in the form of a Fredholm determinant, depending on auxiliary quantum fields. Numerical analysis of correlation functions in the one-dimensional Bose gas with repulsion was carried out in [45] using explicit formulas for form factors obtained within the framework of the ABA.
@@
In this paper, we consider equal-time correlation functions of the one-dimensional bosons with attraction at zero temperature. Unlike dynamical correlations, there is no need to use the form factor expansion in this case. In particular, in [25] a closed representation was obtained for the generating function of the density correlations in any quantum integrable model with a 6-vertex $R$-matrix [^2] A similar representation for the field correlation functions is obtained in this work.. This result is quite applicable to the model of the one-dimensional Bose gas. It should be said, however, that the above-mentioned representation is very cumbersome and contains summation over all possible partitions of the Bethe parameters. In [25], this sum over partitions was reduced to a multiple contour integral of Cauchy type (master-equation [23]), which in turn generates numerous integral representations for correlation functions in the thermodynamic limit [25, 27].
@@ 2 Algebraic Bethe ansatz
A necessary condition for the applicability of the ABA is the existence in the space $\mathcal{H}$ of a special vector called pseudovacuum. In the case of the one-dimensional Bose gas model, this vector coincides with the Fock vacuum $|0\rangle$ (1.3). It is an eigenvector to the diagonal matrix elements of the monodromy matrix and is annihilated by the operator $C(u)$:
@@ 2.3 Composite model
where functions $a^{(j)}(u)$ and $d^{(j)}(u)$ in the one-dimensional Bose gas model are
@@ 2.4 Ground state
To conclude this section, we recall that $C(u)=-B^{\dagger}(u^{*})$ in the attractive one-dimensional Bose gas model. Therefore, we have for the ground state solution
@@
3 Scalar products of Bethe vectors an form factors
@@ 3.1 General scalar products
Equation (3.3) describes the scalar product of two arbitrary off-shell Bethe vectors. However, if at least one of the vectors is on-shell, the sum over the partitions turns into a single determinant of an $N\times N$ matrix [28]. In what follows we will need some particular cases of the scalar product.
@@ 3.2 Density form factor
Representation (3.4) is valid for any integrable model with the $R$-matrix (2.1). In the case of the one-dimensional bosons model, this scalar product is closely related to the form factor of the density operator (1.6) [28]
@@ 3.3 Field form factor
4 Representation for correlation functions
@@
As we noted above, in the case of general solutions of the ordinary and twisted Bethe equations $\bar{u}$ and $\bar{v}$, the sum over partitions (4.7) generates a representation in the form of a multiple Cauchy-type contour integral [23]. However, as we will see below, when calculating the correlation function of densities in the one-dimensional Bose gas with attraction at zero temperature, most of the terms in this sum turn out to be exponentially small in the $\varkappa L\gg 1$ regime.
@@
Below we will need a very particular case of the functions $S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta)$ $S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta)$. We will find an explicit expression for these functions in this particular case.
@@ 5.1 Special values of the parameters $\bar{u}$
$g(u_{k},\bar{v})=\prod_{j=1}^{n}g(u_{k},v_{j}),$ for $\text{for }S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta),$ (5.3)
$g(u_{k},\bar{v})=\prod_{j=1}^{n-1}g(u_{k},v_{j}),$ for $\text{for }S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta).$
@@
Representations similar to (5.1) and (5.2) were obtained in [47]. However, the case $\bar{u}=\bar{\lambda}$ was considered in that work. Therefore, the condition $u_{k+1}=u_{k}+\eta$ was satisfied approximately in the $\varkappa L\gg 1$ regime. The authors of [47] calculated the leading term of the asymptotics, but they did not evaluate the corrections. In this paper, we propose a different method based on recursions (3.11) and (3.16). Using this method, it is easy to calculate the corrections when $\bar{u}=\bar{\lambda}$ in the $\varkappa L\gg 1$ regime.
@@
It is easy to see that under the condition of the proposition
@@
$\mathcal{A}_{n}=\frac{1}{g(u_{n},v_{n-1})}\sum_{k=0}^{+n-2}\frac{\binom{n-2}{k}e^{(k+1)\beta}(-1)^{n-k}}{g(u_{k+2},\bar{v}_{n-1})}-\frac{1}{g(u_{1},v_{n-1})}\sum_{k=0}^{n-2}\frac{\binom{n-2}{k}e^{k\beta}(-1)^{n-k}}{g(u_{k+1},\bar{v}_{n-1})}.$ (5.9)
@@ 5.2 Expansion over $\beta$
In the representations for the correlation functions, we need either the values of the functions $S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta)$ and $S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta)$ at $\beta=0$ or the values of the derivatives of these functions at $\beta=0$. Therefore, we can expand representations (5.1) and (5.2) in a series with respect to the parameter $\beta$ and then restrict ourselves to only the first few terms of the expansion. Then formulas (5.1) and (5.2) are simplified even more.
@@ 5.3 Special values of the parameters $\bar{v}$
$H^{\mathbf{J}}_{n}=n!\sum_{k=0}^{n-1}e^{k\beta}\binom{n-1}{k}\int_{0}^{1}z^{k-s}(1-z)^{n+s-k-1}\,\mathrm{d}z\\ n!\int_{0}^{1}z^{-s}(1-z)^{s}\left(1+z(e^{\beta}-1)\right)^{n-1}\,\mathrm{d}z=\frac{\pi n!}{\sin\pi s}\;{}_{2}F_{1}(1-s,1-n;2;1-e^{\beta}).$ (5.26)
@@ 6 Ground state correlation functions
In this section, we proceed directly to calculating the correlation functions. We will see that the ground state correlation functions for the model of one-dimensional bosons with attraction are described precisely by the special cases of the functions $S^{\mathbf{J}}_{n}(s|\beta)$ and $S^{\mathbf{\Psi}}_{n}(s|\beta)$ given in (5.29).
@@ 6.1 Norm of the ground state
We have mentioned already in the beginning of section 3.2 that the function $S^{\mathbf{J}}_{N}(\bar{v},\bar{u}|\beta)$ describes the scalar product of a twisted dual on-shell Bethe vector $\langle 0|\boldsymbol{C}(\bar{v})$ and ordinary on-shell Bethe vector $\boldsymbol{B}(\bar{u})|0\rangle$. Therefore,
@@ Conclusion
In the case of the one-dimensional Bose gas with repulsion, equations (4.7) and (4.8) ultimately lead to representations for the correlation functions in the form of multiple integrals [25]. The analysis of these integrals is a very difficult task. To date, based on these integral representations, it has been possible to obtain only the asymptotic behavior of correlation functions at a large distance $x$ [27].
However, for the one-dimensional Bose gas with attraction in the $\varkappa L\gg 1$ regime, equations (4.7) and (4.8) are dramatically simplified. Of the entire sum by partitioning of $2N$ roots of the Bethe equations, only $N$ terms make a finite contribution to the correlation. Moreover, these terms are also simplified, since all the determinants included in them are explicitly calculated. As a result, we obtain very simple and explicit results for the correlation functions.
@@
Unfortunately, the method used in this paper is not suitable for computing dynamical correlation functions. Either it should be significantly modified, or resort to form factor decomposition, as was done in [46, 47]. In the latter case, the question arises about summation over the complete set of intermediate states. Recall that in [46, 47], contributions from only two types of excitations were taken into account. It is quite possible that taking into account all types of excitations when calculating dynamic correlations will lead to simpler results.
@@ A.1 Representation for the matrix element in terms of the scalar products
Finally, using (2.25) and again using Bethe equations we arrive at

### v2

@@ 1 Introduction
In this paper, we apply the ABA to the study of correlation functions of a one-dimensional Bose gas (the Lieb–Liniger model) with attractive interaction. The Hamiltonian of this model is given by
@@
The parameter $c$ in (1.1) is a coupling constant. The case of attraction corresponds to $c<0$. Therefore, in what follows we set $c=-\varkappa$ and consider $\varkappa>0$.
The spectral problem for this model was solved in [37, 38]. The ground state energy and the wave function in the attractive case were studied in [39, 40]. A formulation of the model in the ABA language was given in [41, 42]. A series of multiple integrals for correlation functions of this model was derived in [43, 44]. In [29], a representation for the correlation function was obtained in the form of a Fredholm determinant, depending on auxiliary quantum fields. Numerical analysis of correlation functions in the Lieb–Liniger model with repulsion was carried out in [36] using explicit formulas for form factors obtained within the framework of the ABA.
@@
In this paper, we consider equal-time correlation functions of the one-dimensional bosons with attraction at zero temperature. Unlike dynamical correlations, there is no need to use the form factor expansion in this case. In particular, in [25] a closed representation was obtained for the generating function of the density correlations in any quantum integrable model with a 6-vertex $R$-matrix [^2] A similar representation for the field correlation functions is obtained in this work.. This result is quite applicable to the Lieb–Liniger model. It should be said, however, that the above-mentioned representation is very cumbersome and contains summation over all possible partitions of the Bethe parameters. In [25], this sum over partitions was reduced to a multiple contour integral of Cauchy type (master-equation [23]), which in turn generates numerous integral representations for correlation functions in the thermodynamic limit [25, 27].
@@ 2 Algebraic Bethe ansatz
A necessary condition for the applicability of the ABA is the existence in the space $\mathcal{H}$ of a special vector called pseudovacuum. In the case of the Lieb–Liniger model, this vector coincides with the Fock vacuum $|0\rangle$ (1.3). It is an eigenvector to the diagonal matrix elements of the monodromy matrix and is annihilated by the operator $C(u)$:
@@ 2.3 Composite model
where functions $a^{(j)}(u)$ and $d^{(j)}(u)$ in the Lieb–Liniger model are
@@ 2.4 Ground state
To conclude this section, we recall that $C(u)=-B^{\dagger}(u^{*})$ in the attractive Lieb–Liniger model. Therefore, we have for the ground state solution
@@
3 Scalar products of Bethe vectors and form factors
@@ 3.1 General scalar products
Remark 3.1.
Restrictions on the cardinalities of the subsets arise because the highest coefficients exist only for sets of the same cardinality.
Equation (3.3) describes the scalar product of two arbitrary off-shell Bethe vectors. However, if at least one of the vectors is on-shell, the sum over the partitions turns into a single determinant of an $N\times N$ matrix [28]. In what follows we will need some particular cases of the scalar products.
@@ 3.2 Density form factor
Representation (3.4) is valid for any integrable model with the $R$-matrix (2.1). In the case of the Lieb–Liniger model, this scalar product is closely related to the form factor of the density operator (1.6) [28]
@@ 3.3 Field form factor
4 Representations for correlation functions
@@
As we noted above, in the case of general solutions of the ordinary and twisted Bethe equations $\bar{u}$ and $\bar{v}$, the sum over partitions (4.7) generates a representation in the form of a multiple Cauchy-type contour integral [23]. However, as we will see below, when calculating the correlation function of densities in the attractive Lieb–Liniger model at zero temperature, most of the terms in this sum turn out to be exponentially small in the $\varkappa L\gg 1$ regime.
@@
Below we will need a very particular case of the functions $S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta)$ and $S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta)$. We will find an explicit expression for these functions in this particular case.
@@ 5.1 Special values of the parameters $\bar{u}$
$g(u_{k},\bar{v})=\prod_{j=1}^{n}g(u_{k},v_{j}),$ for $\text{for }\quad S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta),$ (5.3)
$g(u_{k},\bar{v})=\prod_{j=1}^{n-1}g(u_{k},v_{j}),$ for $\text{for }\quad S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta).$
@@
Representations similar to (5.1) and (5.2) were obtained in [46]. However, the case $\bar{u}=\bar{\lambda}$ was considered in that work. Therefore, the condition $u_{k+1}=u_{k}+\eta$ was satisfied approximately in the $\varkappa L\gg 1$ regime. The authors of [46] calculated the leading term of the asymptotics. In this paper, we propose a different method based on recursions (3.11) and (3.16).
@@
It is easy to see that under the condition of the proposition,
@@
$\mathcal{A}_{n}=\frac{1}{g(u_{n},v_{n-1})}\sum_{k=0}^{n-2}\frac{\binom{n-2}{k}e^{(k+1)\beta}(-1)^{n-k}}{g(u_{k+2},\bar{v}_{n-1})}-\frac{1}{g(u_{1},v_{n-1})}\sum_{k=0}^{n-2}\frac{\binom{n-2}{k}e^{k\beta}(-1)^{n-k}}{g(u_{k+1},\bar{v}_{n-1})}.$ (5.9)
@@ 5.2 Expansion over $\beta$
In the representations for the correlation functions, we need either the values of the functions $S^{\mathbf{J}}_{n}(\bar{v},\bar{u}|\beta)$ and $S^{\mathbf{\Psi}}_{n}(\bar{v},\bar{u}|\beta)$ at $\beta=0$ or the values of the derivatives of these functions at $\beta=0$. Therefore, we can expand representations (5.1) and (5.2) in series with respect to the parameter $\beta$ and then restrict ourselves to only the first few terms of the expansions. Then formulas (5.1) and (5.2) are simplified even more.
@@ 5.3 Special values of the parameters $\bar{v}$
$H^{\mathbf{J}}_{n}=n!\sum_{k=0}^{n-1}e^{k\beta}\binom{n-1}{k}\int_{0}^{1}z^{k-s}(1-z)^{n+s-k-1}\,\mathrm{d}z\\ =n!\int_{0}^{1}z^{-s}(1-z)^{s}\left(1+z(e^{\beta}-1)\right)^{n-1}\,\mathrm{d}z=\frac{\pi n!}{\sin\pi s}\;{}_{2}F_{1}(1-s,1-n;2;1-e^{\beta}).$ (5.26)
@@ 6 Ground state correlation functions
In this section, we proceed directly to calculating the correlation functions. We will see that the ground state correlation functions for the Lieb–Liniger model with attraction are described precisely by the special cases of the functions $S^{\mathbf{J}}_{n}(s|\beta)$ and $S^{\mathbf{\Psi}}_{n}(s|\beta)$ given in (5.29).
@@ 6.1 Norm of the ground state
We have mentioned already in the beginning of section 3.2 that the function $S^{\mathbf{J}}_{N}(\bar{v},\bar{u}|\beta)$ describes the scalar product of a twisted dual on-shell Bethe vector $\langle 0|\boldsymbol{C}(\bar{v})$ and an ordinary on-shell Bethe vector $\boldsymbol{B}(\bar{u})|0\rangle$. Therefore,
@@ Conclusion
In the case of the Lieb–Liniger model with repulsion, equations (4.7) and (4.8) ultimately lead to representations for the correlation functions in the form of multiple integrals [25]. The analysis of these integrals is a very difficult task. To date, based on these integral representations, it has been possible to obtain only the asymptotic behavior of correlation functions at a large distance $x$ [27].
However, for the attractive Lieb–Liniger model in the $\varkappa L\gg 1$ regime, equations (4.7) and (4.8) are dramatically simplified. Of the entire sum by partitioning of $2N$ roots of the Bethe equations, only $N$ terms make a finite contribution to the correlation. Moreover, these terms are also simplified, since all the determinants included in them are explicitly calculated. As a result, we obtain very simple and explicit results for the correlation functions.
@@
Since representations (4.7), (4.8) are valid for any eigenstate of the Hamiltonian, we can also use them to calculate temperature correlation functions. However, we have to deal with states that contain a mixture of strings of different lengths in this case. Therefore, we no longer have such significant simplifications, despite the fact that string solutions of the Bethe equations are still determined up to exponentially small corrections in the $\varkappa L\gg 1$ regime. Many more terms survive in the sums over partitions, and these terms themselves have a more complex form. This observation further highlights the fact that it is the remarkable ground state structure that is responsible for the dramatic simplifications of the correlation functions at zero temperature in the attractive Lieb–Lineger model. Apparently, this is the main difference between the model under consideration and $1+1$-dimensional models of quantum field theory, in which significant simplifications are also observed in the large system size limit.
Unfortunately, the method used in this paper is not suitable for computing dynamical correlation functions. Either it should be significantly modified, or resort to the form factor decomposition, as was done in [45, 46]. In the latter case, the question arises about summation over the complete set of intermediate states. Recall that in [45, 46], contributions from only two types of excitations were taken into account. It is quite possible that taking into account all types of excitations when calculating dynamic correlations will lead to simpler results.
@@ A.1 Representation for the matrix element in terms of the scalar products
Finally, taking into account (2.25) and using again Bethe equations we arrive at